OpenStudy (anonymous):

Solve & find the general solution of:

OpenStudy (anonymous):

\[\tan^2\frac{ \theta }{ 2 }-1=0\]

OpenStudy (anonymous):

im sorry i can't see anything??

OpenStudy (anonymous):

i didn't get it u cant see <_<

OpenStudy (anonymous):

ii meant i don't see any equation

OpenStudy (anonymous):

well the question itself is an equation

OpenStudy (anonymous):

i think solving for theta gives Theta=2*arctan(1)

OpenStudy (anonymous):

nvm he/shes got it

OpenStudy (anonymous):

\[\tan \frac{ \theta }{ 2 }=\pm \sqrt{\frac{ 1-\cos \theta }{ 1+\cos \theta }}\]

OpenStudy (anonymous):

is that a new question? or was it the first one?

hartnn (hartnn):

why to convert it to cos, keep it in tan tan (x/2) = +1 or -1

OpenStudy (anonymous):

what how?

hartnn (hartnn):

tan^2 (x/2) -1 =0 tan^2 (x/2) = 1 tan (x/2) =1 or tan (x/2) = -1

hartnn (hartnn):

x/2 = 45 (or pi/4) or x/2 = -(pi/4) x=...? or .... ?

OpenStudy (anonymous):

tan x/2 =1 how to do it?

hartnn (hartnn):

tan value = 1 at angles, 45, 45+180,45+2*180,... 45 +n*180

hartnn (hartnn):

so, tan (x/2) = 1 implies (x/2) = pi/4 +n * pi

hartnn (hartnn):

x= 2*pi/4 + 2*n*pi

hartnn (hartnn):

x= pi/2 +2n *pi same for tan (x/2) =-1

OpenStudy (anonymous):

\[{n \pi \pm \frac{ \pi }{ 2 }}\]

OpenStudy (anonymous):

this is the answer

hartnn (hartnn):

and we got the same ? its like tan (x/2) = 1 implies (x/2) = pi/4 x= 2*pi/4 = pi/2 and since tan function is periodic in pi, x = pi/2 +n*pi for tan (x/2) = -1 x/2 = -pi/4 x = -pi/2 +n*pi

hartnn (hartnn):

got that ^ ?

OpenStudy (anonymous):

ok